∫cos 5 x√sinxdx

∫cos5x√sinxdx       

Grade:12

1 Answers

Ashwin Muralidharan IIT Madras
12 years ago

Hi Suhas,

 

MAke the sub sinx = t^2, so that cosxdx = 2tdt.

And cos^4x = (1-sin^2 x)^2 = (1-t^4)^2

So that the integral becomes ∫[(1-t^4)^2]2t^2dt.

Open the brackets. Directly integrable.

 

Hope it helps.

 

Best Regards,

Ashwin (IIT MadraS).